# VF-Plan: Bridging the Art Gallery Problem and Static LiDAR Scanning with Visibility Field Optimization

### v1

@@ Abstract
Viewpoint planning is crucial for 3D data collection and autonomous navigation, yet existing methods often miss key optimization objectives for static LiDAR, resulting in suboptimal network designs. The Viewpoint Planning Problem (VPP), which builds upon the Art Gallery Problem (AGP), requires not only full coverage but also robust registrability and connectivity under limited sensor views. We introduce a greedy optimization algorithm that tackles these VPP and AGP challenges through a novel Visibility Field (VF) approach. The VF captures visibility characteristics unique to static LiDAR, enabling a reduction from 2D to 1D by focusing on medial axis and joints. This leads to a minimal, fully connected viewpoint network with comprehensive coverage and minimal redundancy. Experiments across diverse environments show that our method achieves high efficiency and scalability, matching or surpassing expert designs. Compared to state-of-the-art methods, our approach achieves comparable viewpoint counts (VC) while reducing Weighted Average Path Length (WAPL) by approximately 95%, indicating a much more compact and connected network. Dataset and source code will be released upon acceptance.
keywords:
Viewpoint Optimization , Terrestrial Laser Scaning , Visibility Field , Medial axis transform , Greedy searching , BIM
@@
[label1]organization=School of Computer Science and Artificial Intelligence, Wuhan University of Technology,addressline=122 Luoshi Rd, city=Wuhan, postcode=430070, state=Hubei, country=China
@@
[label2]organization=College of Computer Science and Technology, Zhejiang University, addressline=866 Yuhangtang Rd, city=Hangzhou, postcode=310058, state=Zhejiang, country=China
[…]
[label3]organization=The Advanced Laser Technology Laboratory of Anhui Province, addressline= 850 Huangshan Rd, city=Hefei, postcode=230031, state=Anhui, country=China
{graphicalabstract}
Viewpoint optimization for static LiDAR Scanning. (a) The input floor plan; (b) Optimized viewpoints indicated by green dots, with the Visibility Field displayed as the background image; (c) and (d) Simulated LiDAR points from optimal viewpoints, shown in 2D and 3D respectively, with each scan distinguished by a unique color.
{highlights}
Unified Solution to VPP and AGP: Our method addresses the combined challenges of viewpoint planning and the Art Gallery Problem, achieving efficient coverage with minimal viewpoints and robust network connectivity.
Visibility Field for Dimensionality Reduction: We introduce a continuous Visibility Field tailored to static LiDAR systems, reducing the optimization space from 2D to 1D by leveraging structural elements like medial axis and joints.
Greedy Algorithm for Efficient Coverage: Our greedy Algorithm utilizes the VF’s structural insights to construct a minimal, fully connected viewpoint network, ensuring both robust coverage and network connectivity.
@@ 1 Introduction
The Viewpoint Planning Problem (VPP) has emerged as a critical research spot due to the growing use of LiDAR and camera sensors in applications such as 3D reconstruction, autonomous navigation, and Scan-to-BIM applications [41, 30]. VPP focuses on determining the optimal placements and configurations of sensors to achieve specific reconstruction goals, such as maximizing observation completeness and minimizing occlusions, while balancing efficiency and data quality. Both model-free and model-based algorithms for viewpoint and path planning have gained significant attention due to their applicability across various industries, including robotics, construction, and surveillance [31].
Conceptually, VPP is rooted in the Art Gallery Problem (AGP), a classic challenge in computational geometry that seeks the minimal number of “guards” required to cover a given area, typically a polygon, with full 360° visibility [25]. AGP is often simplified by constraining guard placements to polygon vertices; however, this simplification does not capture the true computational complexity of the problem. In its generalized form, AGP is $\exists\mathbb{R}$-complete, meaning it is at least as hard as any NP-complete problem, with no known efficient solutions [11, 1]. This NP-hard nature has inspired substantial research interest, particularly in constrained versions that make idealized assumptions, such as complete visibility and static scenes [11, 41].
Figure 1: Viewpoint optimization for static LiDAR Scanning. (a) The input floor plan; (b) Optimized viewpoints indicated by green dots, with the Visibility Field displayed as the background image; (c) and (d) Simulated LiDAR points from optimal viewpoints, shown in 2D and 3D respectively, with each scan distinguished by a unique color.
VPP, however, introduces additional complexities beyond those of AGP. Unlike static guards, sensors in VPP must maintain overlapping fields of view for accurate data registration, accommodate limited and directional visibility ranges, and meet high accuracy requirements for 3D scene reconstruction. These real-world demands, especially in cluttered or partially obstructed environments, render VPP a more challenging NP-hard problem, requiring innovative optimization strategies beyond those used for AGP [41, 1].
Our work addresses these challenges by introducing a novel greedy approach based on the concept of a Visibility Field (VF), specifically designed for static LiDAR applications. The VF captures the unique visibility characteristics of LiDAR, including constraints on range and incident angles, enabling a significant reduction in computational complexity. By focusing on critical points such as medial axis and joints within the VF, we effectively reduce the optimization space from 2D to 1D, facilitating the construction of a minimal, fully connected viewpoint network that achieves comprehensive coverage with reduced redundancy. Our main contributions are as follows:
@@
    Unified Solution to VPP and AGP: Our method addresses the combined challenges of viewpoint planning and the Art Gallery Problem, achieving efficient coverage with minimal viewpoints and robust network connectivity.
@@
    Visibility Field for Dimensionality Reduction: We introduce a continuous Visibility Field tailored to static LiDAR systems, reducing the optimization space from 2D to 1D by leveraging structural elements like medial axis and joints.
@@
    Greedy Algorithm for Efficient Coverage: Our greedy algorithm utilizes the VF’s structural insights to construct a minimal, fully connected viewpoint network, ensuring both robust coverage and network connectivity.
@@ 2.1 Station-Based Scanning
Sensor network design shares similarities with the Planning for Scanning (P4S) problem, observed in applications like surveillance cameras [32, 55, 16], directional sensors [15], 5G base stations [42, 45], stealth game [49], and station-based LiDAR scanning [2]. For LiDAR, the goal is to cover the entire scene with comprehensive wall scanning while ensuring data quality criteria such as completeness and accuracy [29, 2].
Advancements in LiDAR technology have led to P4S methods using prior information like 2D drawings and 3D models [29]. These methods have been applied to civil infrastructures [50], space frame structures [29], rebar detection [27, 28], and landslide monitoring [54]. In cases where prior information is outdated or unavailable, low-resolution preliminary scans are required, either indoors [39] or outdoors using UAVs [26].
P4S optimizations aim to maximize coverage and overlap while minimizing the number of viewpoints [19, 7]. Methods like CMA Evolution Strategy [40] and regular grid sampling [21, 33] increase the likelihood of finding optimal viewpoints but also raise computational complexity. Jia’s approach using grid-based sampling [20, 21, 22] scores candidate viewpoints based on observed wall segments, though this approach is limited by its reliance on reference objects for registration. Noichl [35] improve Jia’s method into 3D scene for an uniform covering by triangulating 3D scene into meshes and checking their visibility. For scenarios lacking BIM priors, structural elements can be surveyed and detected using low-cost sensors to serve as a preliminary prior for static station planning. A precise scan is then performed in a stop-and-go manner using a static scanner mounted on a robotic dog [9, 18].
Our approach addresses these limitations by using a continuous Visibility Field (VF) model, which reduces optimization from 2D to 1D by focusing on structural convergence points like medial axis and joints, thus minimizing computation and enhancing adaptability.
@@ 2.2 Mobile-Based Exploration
Coverage Path Planning (CPP) focuses on computing optimal, collision-free paths for robots to fully cover target areas [44]. Techniques like Rapidly-exploring Random Trees (RRT) and their variants (RRT* [23], RRT# [37]) are widely used for autonomous exploration in unknown spaces, including UAVs [31] and wheeled robots [38]. RRT-based methods maximize information gain while minimizing travel distance [17, 47, 53].
@@
Methods utilizing tensor fields for guiding robot movements enable efficient path routing in partially reconstructed indoor scenes [48, 46]. Recent approaches like the Signed Distance Field (SDF) and Hamilton-Jacobi skeleton further improve path planning in known environments [36]. These concepts share similarities with our visibility field, which leverages converging lines for efficient viewpoint planning.
@@
For known environments, algorithms like the receding horizon ”next-best-view” approach [3] and dual-resolution mapping schemes [6] balance detailed local mapping with efficient global exploration. For applications in autonomous driving and exploration, recent advances such as ”LookOut” [12] propose diverse multi-future prediction models that enhance autonomous navigation by predicting possible trajectories, balancing safety, and efficiency.
3 Theoretical Foundations and Global Optimality
This section formulates the Viewpoint Planning Problem (VPP) for static LiDAR scanning in a 2D polygonal environment and establishes the theoretical basis for restricting candidate viewpoints to the medial axis (MA). We first present the problem definition, including coverage and connectivity requirements, and then show that any line segment visible from an interior point is also visible from a corresponding set of MA points. This demonstrates that restricting the viewpoint search space to the MA incurs no loss of coverage capability. Finally, we prove that a minimal solution confined to the MA is globally optimal.
3.1 Problem Definition
Let $P$ be a 2D polygonal region (such as a floor plan) with possible obstacles or holes, while remaining topologically connected. The polygon’s boundary (including internal obstacles) is discretized into $n$ small line segments:
$L=\{\,l_{1},\;l_{2},\;\dots,\;l_{n}\},$
which represent the target structures (e.g., walls or windows) to be covered by static LiDAR scanners (hereafter referred to as viewpoints).
Candidate Viewpoints
We aim to position a set of viewpoints such that all segments in $L$ are fully visible, while the chosen viewpoints remain connected under an overlap criterion. Let $V=\{\,v_{1},\;v_{2},\;\dots,\;v_{m}\}$ be the set of candidate viewpoint locations. We will later establish that it is sufficient to confine $V$ to the medial axis of $P$, reducing a 2D search to a simpler 1D structure.
Visibility
we define a Completeness Table $C$, where $C_{ij}$ indicates whether viewpoint $V_{i}$ observes line segment $L_{j}$
$C_{ij}=\begin{cases}1&\text{if }v_{i}\text{ has an unobstructed view of }l_{j}\\ 0&\text{otherwise}\end{cases}$ (1)
Connectivity via Overlap
Consider any pair of viewpoints $(v_{i},v_{k})$. Let $O_{ik}$ be their overlap ratio, i.e., the fraction of segments in $L$ that both $v_{i}$ and $v_{k}$ can simultaneously observe. For a threshold $\tau$, an edge $(v_{i},v_{k})$ exists in an overlap graph $G=(V,E)$ if $O_{ik}\geq\tau$. A chosen subset $S\subseteq V$ must then form a connected subgraph under this threshold.
Mathematical Formulation
We seek a minimal subset $S\subseteq V$ of viewpoints such that:
  1. (a)
    Coverage: Each segment $l_{j}\in L$ is observed by at least one viewpoint in $S$, i.e.,
    $\forall\,l_{j}\in L,\;\exists\,v_{i}\in S:\;C_{ij}=1.$ (2)
  2. (b)
@@ 3.1 Problem Definition
    Connectivity: The induced subgraph $(S,E_{S})$ is connected under the overlap threshold $\tau$, where
@@
The optimization problem becomes:
@@
Since problem (4) generalizes the set cover problem (with additional connectivity constraints), it remains NP-hard. Nonetheless, by restricting $V$ to the medial axis (MA), we retain optimality while reducing the search complexity considerably.
3.2 Skeleton Completeness and Visibility Propagation
Medial Axis (MA)
The MA of $P$ is the set of points inside $P$ that are equidistant to at least two boundary edges. In Blum’s grassfire transform [4], if the boundary of $P$ is ignited simultaneously, the inward-propagating flames collide and form the locus of the MA.
Theorem 1 (Skeleton Completeness).
Suppose an interior point $p\in P$ sees a set of boundary segments $L_{p}\subseteq L$. Then there exists a set of points $\{q_{l}\}\subseteq\mathrm{MA}(P)$ such that
$L_{p}\;\;\subseteq\;\;\bigcup_{l\in L_{p}}\mathrm{Vis}(q_{l}).$
Hence, for any interior coverage provided by $p$, there exists an MA-based set of points whose collective coverage reproduces that of $p$.
Sketch of Proof. Using Blum’s “fire transformation,” the visibility from every segment propagates inward until it intersects the medial axis $\mathrm{MA}(P)$. Each line segment $l_{j}\in L_{p}$ must intersect $\mathrm{MA}(P)$ at some point $q_{j}$ visible to $l_{j}$. Thus,
$\mathrm{Vis}(p)\;\;\subseteq\;\;\bigcup_{l_{j}\in\mathrm{Vis}(p)}\,\mathrm{Vis}(q_{j}).$
Consequently, although a single point on the MA may not match $p$’s entire coverage, a suitable collection of MA points can, illustrated in 2. Restricting $V$ to $\mathrm{MA}(P)$ therefore does not diminish the achievable coverage.
3.3 Global Minimal Cover–Connectivity on the Skeleton
By Theorem 1, any coverage by interior points can be equivalently transferred to a suitable set of points on the medial axis. Hence, in the formulation (4), restricting $V$ to $\mathrm{MA}(P)$ (including any branch or joint nodes) captures all feasible coverage solutions.
Theorem 2 (Global Minimal Cover–Connectivity).
Let $S^{*}\subseteq\mathrm{MA}(P)$ be a minimal solution to (4) that covers all $l_{j}\in L$ and satisfies the overlap-based connectivity requirement. Then $S^{*}$ is irreducible (i.e., removing any $v_{i}\in S^{*}$ breaks coverage or connectivity), and no smaller interior configuration can outperform $S^{*}$. Therefore, $S^{*}$ is globally optimal.
Implication and Heuristic Search
Exploiting the symmetry of the medial axis, no point outside $\mathrm{MA}(P)$ can provide superior coverage or connectivity. Consequently, a minimal cover–connectivity solution restricted to the skeleton cannot be improved by adding non-skeleton points. Hence, solving (4) directly on $\mathrm{MA}(P)$ yields a globally optimal arrangement in $P$. Although the problem remains NP-hard, standard polynomial-time greedy algorithm (e.g., [10]) often produce near-optimal results in practice. Once obtained, the minimal MA-based solution is irreducible, further affirming its global optimality for the entire polygon.
Figure 3: Overview of the Viewpoint Network (VPN) optimization based on the Visibility Field (VF). (a) Input floor plan; (b) Visibility Field (VF); (c) Distance field of VF; (d) Converging lines and points; (e) Optimized Viewpoint Network.
@@ 4.1 Overview
Our proposed VF-Plan addresses the NP-hard complexity of the VPP through a greedy approach centered on the VF. Given a structural floor plan, VF-Plan constructs a Viewpoint Network (VPN) that achieves comprehensive coverage with minimal viewpoints, ensuring connectivity and registrability. Each viewpoint is selected to form a fully connected observation network, minimizing registration errors via Least Squares Fitting (LSF).
As illustrated in Figure 3, the VF-Plan pipeline starts from a floor plan with structural elements (walls, columns, doors, windows) and proceeds through the steps of VF computation, distance field generation, identification of medial axis and joint points, and initialization of candidate viewpoints. A greedy algorithm then refines these candidates into a minimal, connected VPN. To optimize the computationally intensive task of determining visible wall segments from each viewpoint, we employ a BSP-tree[14] to accelerate processing.
4.2 Visibility Field
Laser scanning is limited to visible surfaces, leaving occluded areas unobserved. Therefore, visibility and coverage analysis are essential for viewpoint planning, with key objectives of maximizing observation completeness and registrability. The field of view is central to VPN optimization, defining the information captured at each viewpoint and guiding the selection of optimal viewpoints.
A common method for evaluating the information scanned by a LiDAR sensor is based on the length of observed walls [22]. Typically, wall length is counted either by ensuring both endpoints are visible or by partitioning walls into uniformly sampled segments within the scanner’s range. For 3D cases, visibility is evaluated by counting the number of observed voxels. However, these approaches lack normalization, yielding values that vary with discretization resolution and do not account for differences in distance or orientation to the scanner, impacting computational stability and load without improving coverage accuracy.
Our approach instead uses the valid scanned angle to integrate relevant factors into a single index, uniformly scaled from 0 to $2\pi$. We define a static LiDAR scanner with a 360° field of view in the X-Y plane, a range from $r_{\text{min}}$ to $r_{\text{max}}$, and uniform scanning speed, as shown in Figure 4. Only objects within this range are observable. Generally, closer viewpoints collect more information, but excessive proximity can occlude other parts of the field. Scanning is performed in discrete observation angles, with observed information defined as the valid observed angle, ensuring alignment with scanner properties.

### v2

@@ Abstract
Viewpoint planning is critical for efficient 3D data acquisition in applications such as 3D reconstruction, building life-cycle management, navigation, and interior decoration. However, existing methods often neglect key optimization objectives specific to static LiDAR systems, resulting in redundant or disconnected viewpoint networks. The viewpoint planning problem (VPP) extends the classical Art Gallery Problem (AGP) by requiring full coverage, strong registrability, and coherent network connectivity under constrained sensor capabilities. To address these challenges, we introduce a novel Visibility Field (VF) that accurately captures the directional and range-dependent visibility properties of static LiDAR scanners. We further observe that visibility information naturally converges onto a 1D skeleton embedded in the 2D space, enabling significant searching space reduction. Leveraging these insights, we develop a greedy optimization algorithm tailored to the VPP, which constructs a minimal yet fully connected Viewpoint Network (VPN) with low redundancy. Experimental evaluations across diverse indoor and outdoor scenarios confirm the scalability and robustness of our method. Compared to expert-designed VPNs and existing state-of-the-art approaches, our algorithm achieves comparable or fewer viewpoints while significantly enhancing connectivity. In particular, it reduces the weighted average path length by approximately 95%, demonstrating substantial improvements in compactness and structural efficiency. Code is available at https://github.com/xiongbiaostar/VFPlan.
{CCSXML}
<ccs2012> <concept> <concept_id>10003752.10010061.10010063</concept_id> <concept_desc>Theory of computation Computational geometry</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
@@
[500]Theory of computation Computational geometry
@@ 1 Introduction
The VPP has emerged as a critical research spot due to the growing use of LiDAR and camera sensors in applications such as 3D reconstruction, autonomous navigation, facility assessment, building life-cycle management, and interior decoration applications [SRR03, MLL23]. VPP focuses on determining the optimal placements and configurations of sensors to achieve specific reconstruction goals, such as maximizing observation coverage and minimizing occlusions, while balancing efficiency and data quality. Both model-free and model-based algorithms for viewpoint and path planning have gained significant attention due to their applicability across various industries, including robotics, construction, and surveillance [MHS23].
Conceptually, VPP is rooted in the AGP, a classic challenge in computational geometry that seeks the minimal number of “guards” required to cover a given area, typically a polygon, with full 360° visibility [LL86]. AGP is often simplified by constraining guard placements to polygon vertices; however, this simplification does not capture the true computational complexity of the problem. In its generalized form, AGP is $\exists\mathbb{R}$-complete, meaning it is at least as hard as any NP-complete problem, with no known efficient solutions [CdRdS11, AAM21]. This NP-hard nature has inspired substantial research interest, particularly in constrained versions that make idealized assumptions, such as complete visibility and static scenes [CdRdS11, SRR03].
VPP, however, introduces additional complexities beyond those of AGP. Unlike static guards, sensors in VPP must maintain overlapping fields of view for accurate data registration, accommodate limited and directional visibility ranges, and meet high accuracy requirements for 3D scene reconstruction. These real-world demands, especially in cluttered or partially obstructed environments, render VPP a more challenging NP-hard problem, requiring innovative optimization strategies beyond those used for AGP [SRR03, AAM21].
We address the challenges of static LiDAR coverage planning by introducing a novel greedy algorithm based on the concept of a Visibility Field (VF). The VF encapsulates the physical constraints of LiDAR sensors, including limited sensing range and angle-dependent visibility, enabling a substantial reduction in computational complexity. We formally demonstrate that visibility information is inherently concentrated along a one-dimensional skeleton embedded within the two-dimensional space. This observation allows us to restrict the search space from 2D to 1D by targeting critical structural features such as the medial axis and junctions within the VF. Consequently, we construct a minimal, fully connected network of viewpoints that achieves complete scene coverage with reduced redundancy. Our main contributions are summarized as follows:
@@
    Unified Solution to VPP and AGP: Our method addresses the combined challenges of viewpoint planning and the AGP, achieving efficient coverage with minimal viewpoints and robust network connectivity.
@@
    Visibility Field for Dimensionality Reduction: We introduce a continuousVF tailored to static LiDAR systems, reducing the optimization space from 2D to 1D by leveraging structural elements like medial axis and joints.
@@
    Greedy Algorithm for Efficient Coverage: Our greedy algorithm utilizes the VF’s structural insights to construct a minimal, fully connected VPN, ensuring both robust coverage and network connectivity.
Figure 1: Pipeline of viewpoint network (VPN) optimization using VF. (a) Input floor plan with walls, doors, and windows; (b) Computed VF encoding the valid observed angle at each pixel; (c) Distance field derived from the VF; (d) Extraction of converging lines (medial axis) and converging points (joints) used as candidate viewpoints; (e) Final optimized VPN ensuring full coverage and strong connectivity.
@@ 2.1 Station-Based Scanning
Sensor network design shares similarities with the Planning for Scanning (P4S) problem, observed in applications like surveillance cameras [MADR11, ZYCH13, GCW15], directional sensors [FG09], 5G base stations [SSZ20, WZL20], stealth game [XTV14], and station-based LiDAR scanning [ABT21]. For LiDAR, the goal is to cover the entire scene with comprehensive wall scanning while ensuring data quality criteria such as completeness and accuracy [LXHL21, ABT21].
Advancements in LiDAR technology have led to P4S methods using prior information like 2D drawings and 3D models [LXHL21]. These methods have been applied to civil infrastructures [XWYZ25], space frame structures [LXHL21], rebar detection [LKL21, LKS25], and landslide monitoring [ZCH23]. In cases where prior information is outdated or unavailable, low-resolution preliminary scans are required, either indoors [QWTW21] or outdoors using UAVs [LLZ22].
P4S optimizations aim to maximize coverage and overlap while minimizing the number of viewpoints [IGFER22, CSW25]. Methods like CMA Evolution Strategy [RLGA22] and regular grid sampling [JL19, NB22] increase the likelihood of finding optimal viewpoints but also raise computational complexity. Jia’s approach using grid-based sampling [JL18, JL19, JL22] scores candidate viewpoints based on observed wall segments, though this approach is limited by its reliance on reference objects for registration. Noichl [NLB24] improve Jia’s method into 3D scene for an uniform covering by triangulating 3D scene into meshes and checking their visibility. For scenarios lacking BIM priors, structural elements can be surveyed and detected using low-cost sensors to serve as a preliminary prior for static station planning. A precise scan is then performed in a stop-and-go manner using a static scanner mounted on a robotic dog [CKP25, HG25].
Our approach addresses these limitations by using a continuous VF model, which reduces optimization from 2D to 1D by focusing on structural convergence points like medial axis and joints, thus minimizing computation and enhancing adaptability.
@@ 2.2 Mobile-Based Exploration
Coverage Path Planning (CPP) focuses on computing optimal, collision-free paths for robots to fully cover target areas [TMMA21]. Techniques like Rapidly-exploring Random Trees (RRT) and their variants (RRT* [KF11], RRT# [OF15]) are widely used for autonomous exploration in unknown spaces, including UAVs [MHS23] and wheeled robots [PSC23]. RRT-based methods maximize information gain while minimizing travel distance [HNH18, XYH18, ZYX21].
@@
Methods utilizing tensor fields for guiding robot movements enable efficient path routing in partially reconstructed indoor scenes [XZY17, XZD24]. Recent approaches like the Signed Distance Field (SDF) and Hamilton-Jacobi skeleton further improve path planning in known environments [NLMC23]. These concepts share similarities with our VF, which leverages converging lines for efficient viewpoint planning.
@@
For known environments, algorithms like the receding horizon "next-best-view" approach [BKA16] and dual-resolution mapping schemes [CZR23] balance detailed local mapping with efficient global exploration. For applications in autonomous driving and exploration, recent advances such as "LookOut" [CCS21] propose diverse multi-future prediction models that enhance autonomous navigation by predicting possible trajectories, balancing safety, and efficiency.
3 Problem Definition and Global Optimality
We formalize the VPP for static LiDAR scanning across a collection of two-dimensional polygonal regions $\mathcal{P}=\{P_{1},P_{2},\dots,P_{k}\}$, each comprising structural boundaries and possible internal obstacles. The boundaries of these regions are uniformly discretized into a set of equal-length line segments $L=\{l_{1},l_{2},\dots,l_{n}\}$, representing the structural elements to be scanned. The objective is to select a set of viewpoints $V=\{v_{1},v_{2},\dots,v_{m}\}$ that satisfies two core constraints: coverage and connectivity.
Coverage: Every segment $l_{j}\in L$ must be visible from at least one viewpoint $v_{i}\in V$. We formalize this visibility relationship using a binary coverage table $C$, where each entry $C_{ij}$ indicates whether $v_{i}$ observes $l_{j}$:
$C_{ij}=\begin{cases}1&\text{if }v_{i}\text{ sees }l_{j},\\ 0&\text{otherwise}.\end{cases}$ (1)
@@ 3 Problem Definition and Global Optimality
Connectivity: Let the overlap ratio $O_{ik}$ as the fraction of $L$ observed by both viewpoints $v_{i}$ and $v_{k}$. The induced subgraph $(S,E_{s})$ is connected under the overlap threshold $\tau$, where
@@
We aim to find the smallest subset $S\subseteq V$ such that:
@@
This NP-hard problem extends the classical set cover problem by incorporating additional connectivity constraints. We demonstrate that restricting $V$ to the medial axis (MA) preserves both complete coverage and network connectivity. As observation data in the VF naturally converge toward the MA and its joint points, placing viewpoints along these skeletal structures ensures comprehensive coverage. This property reduces VPN optimization from a 2D search to a 1D problem constrained to the MA. Consequently, a minimal cover on the MA remains globally optimal. Formal proofs of visibility completeness and global optimality for this reduction are provided in Appendix A.
@@ 4.1 Overview
Our proposed method, VF-Plan, addresses the NP-hard complexity of the View Planning Problem (VPP) through a greedy strategy centered on the VF. Given a structural 2D floorplan, VF-Plan constructs a VPN that ensures full coverage with a minimal number of viewpoints, while maintaining connectivity and registrability. The selected viewpoints form a fully connected observation network, which can subsequently be used to minimize registration errors via Least Squares Fitting (LSF).
As shown in fig.˜1, VF-Plan starts from a floorplan containing structural elements such as walls, columns, doors, and windows. The process includes VF computation, medial axis and joint-point extraction, and candidate viewpoint initialization. Overlap ratios between candidates are then evaluated, and a greedy algorithm selects the minimal subset forming a connected VPN. To accelerate the visibility checks for each viewpoint, we employ a BSP-tree [FKN80] with balanced splitting and no depth limit, starting from the median line at each iteration. Prior to BSP-tree construction and VF computation, the floorplan is preprocessed into closed polygons while preserving all structural elements.
4.2 Visibility Field (VF)
Laser scanning is limited to visible surfaces, leaving occluded areas unobserved. Therefore, visibility and coverage analysis are essential for viewpoint planning, with key objectives of maximizing observation Coverage and registrability. The field of view is central to VPN optimization, defining the information captured at each viewpoint and guiding the selection of optimal viewpoints.
A common method for evaluating the information scanned by a LiDAR sensor is based on the length of observed walls [JL22]. Typically, wall length is counted either by ensuring both endpoints are visible or by partitioning walls into uniformly sampled segments within the scanner’s range. For 3D cases, visibility is evaluated by counting the number of observed voxels. However, these approaches lack normalization, yielding values that vary with discretization resolution and do not account for differences in distance or orientation to the scanner, impacting computational stability and load without improving coverage accuracy.
Our approach instead uses the valid scanned angle to integrate relevant factors into a single index, uniformly scaled from 0 to $2\pi$. We define a static LiDAR scanner with a 360° field of view in the X-Y plane, a range from $r_{\text{min}}$ to $r_{\text{max}}$, and uniform scanning speed, as shown in fig.˜2. Only objects within this range are observable. Generally, closer viewpoints collect more information, but excessive proximity can occlude other parts of the field. Scanning is performed in discrete observation angles, with observed information defined as the valid observed angle, ensuring alignment with scanner properties.